OpenStudy (anonymous):

Complete the square for the expression. Write the resulting expression as a binomial squared. x2 - 12x + ____

OpenStudy (anonymous):

hmmm.. ok do you know the formula for the expansion of \[(a-b)^{2}\] ?

OpenStudy (anonymous):

x^2 - 12x + 36 = (x + 6)^2 x^2 - 12x + 144 = (x - 12)^2 x^2 - 12x + 36 = (x - 6)^2 x^2 - 12x + 144 = (x + 12)^2 which one is it?

OpenStudy (anonymous):

take the term in front of x, divide it by 2 and square it to get the first part. (-12/2)^2=36 x^2-12x+36

OpenStudy (anonymous):

x^2 - 12x + 36 = (x - 6)^2 or x^2 - 12x + 36 = (x + 6)^2

OpenStudy (anonymous):

Since the 12 is negative it would be (x-6)^2